OpenStudy (anonymous):

Given the following system of equations, identify the type of system. x + y = 6 y = 3 - x consistent inconsistent equivalent b

OpenStudy (anonymous):

solve for y in both equations.

OpenStudy (anonymous):

kk one sec

OpenStudy (anonymous):

x + y = 6 solve for y y = 3 - x already solved for y y=-x+3 rearanged the above one. Basicly get it in y=mx + b mode. you got 3 cases now 1. if ms are different you got consistent 2. A. If m same and b same you got equivlient B. if m same and b different then Inconsistent.

OpenStudy (anonymous):

-3 and 9 ??

OpenStudy (anonymous):

x + y = 6 y=6-x ------ y=-x+6 y=-x+3 Inconsistant

OpenStudy (anonymous):

same m of -1 yet 2 different bs make it inconsistent.

OpenStudy (anonymous):

so inconsistant ?

OpenStudy (anonymous):

Yes inconsistent/ no solution